Electric Potential Energy of Three Point Charges

Three equal point charges, each with charge 1.65 , are placed at the vertices of an equilateral triangle whose sides are of length 0.500 . What is the electricpotential energy of the system? (Take as zero the potential energy of the three charges when they are infinitely far apart.)
Use = 8.85×10-12 for the permittivity of free space.
Answer #1

potential energy = kq^2/r + 2*kq^2/r =3kq^2/r
so,P.E = 3*8.85*10^-12*1.65^2*10^-12/.5 = 1.45*10^-22 J


I took the magnitude of charge to be 1.65μC (= 1.65*10^-6C) and the radius = 0.5m

answered by: S.Dunn